# Korean Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

### v2

@@ Abstract
Investigations into Aspect-Based Sentiment Analysis (ABSA) for Korean restaurant reviews are notably lacking in the existing literature. Our research proposes an intuitive and effective framework for ABSA in low-resource languages such as Korean. It optimizes prediction labels by integrating translated benchmark and unlabeled Korean data. Using a model fine-tuned on translated data, we pseudo-labeled the actual Korean NLI set. Subsequently, we applied LaBSE and MSP-based filtering to this pseudo-NLI set as implicit feature, enhancing Aspect Category Detection and Polarity determination through additional training. Incorporating dual filtering, this model bridged dataset gaps, achieving positive results in Korean ABSA with minimal resources. Through additional data injection pipelines, our approach aims to utilize high-resource data and construct effective models within communities, whether corporate or individual, in low-resource language countries. Compared to English ABSA, our framework showed an approximately 3% difference in F1 scores and accuracy. We release the dataset and our code for Korean ABSA, at this link [^1] https://github.com/namkibeom/KPC-cF.
@@ 2.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur and Turchi Balahur and Turchi (2012) has shown that Machine Translation (MT) systems can obtain training data for languages other than English in general sentiment classification. However, existing research using translation data for alignment and alignment-free methods Li et al. (2020); Zhang et al. (2021) inadequately address the challenge of universal knowledge transfer for linguistic subgroups like Korean. Also, although it was a different domain at Zhou et al. Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature $\Phi$ between translated source data $\bm{D_{S}}$ and actual target language data $\bm{D_{T}}$. Therefore, we attempted the following two phases to assess the impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Fig. 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, similar to the findings of Hoshino et al. (2024), the optimal baseline model for Korean reasoning is identified from the pool of models trained and assessed utilizing the translation dataset. In Phase 2, we fine-tune the baseline model $\Psi_{pre}(\Phi(D_{S});\theta_{D_{S}})$, which was effective in training on $D_{S}$, by additionally incorporating pseudo-labeled actual $D_{T}$. Employing the tuned model $\Psi_{post}(\Phi(D_{T});\theta_{{D_{S}}\rightarrow D_{T}})$, we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, LaBSE and confidence score filtering are performed to enhance implicit features $\Phi(D_{T})$. Detailed description of our Language Adaptation for aligned ABSA task is provided in Appendix B.
@@
7:         $score_{MSP}\leftarrow$ $\textit{MSP}_{\Psi_{pre}}$($y\mid x_{s},x_{a}$)
@@ 2.2 LaBSE based Filtering
We aimed to apply the LaBSE scoring to the actual Korean data $D_{T}$, KR3, intending to facilitate flexible learning compared to the translated data $D_{S}$, Kor-SemEval (see Appendix C.1).
@@ 2.3 Confidence score Filtering
Meanwhile, we need to develop a classifier $\Psi_{post}$ capable of optimal predictions on the $D_{T}$ test, which can be considered as out-of-distribution data separate from the $D_{S}$. Drawing on previous research Arora et al. (2021), we expect that language shifts ($D_{S}\leftrightarrow D_{T}$) embody both Background and Semantic shift characteristics. To ensure robust learning in both aspect detection and sentiment classification, we introduce additional thresholding on Maximum Softmax Probability (MSP; Hendrycks and Gimpel 2017) after LaBSE-based filtering on the $D_{T}$ train set. When considering an input $x=(x_{s},x_{a})\in\mathcal{X}$ and its corresponding pseudo-label $y\in\mathcal{Y}$, the score $s(x)$ for MSP is expressed as:
$s_{\textit{MSP}}(x)=\max_{k\in\mathcal{Y}}p_{\text{model}}(y=k\mid x).$ (4)
@@
Table 2: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to total number (i.e., equivalent to Kor-SemEval) and capacity of samples from KR3 Train. The performances where filtered models outperform non-filtered ones under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Appendix B.3.
@@ 3.1 Main Experiment: KR3 Test Set
Based on the results from Kor-SemEval, we observed analogous patterns between mBERT and XLM-RBase, notwithstanding their distinct properties (see Appendix D). Accordingly, We opted for the Baseline-NLI approach (i.e., BaselinemBERT, XLM-R), which demonstrated the best performance, as the base model for Phase 2.
@@
Results on the KR3 test set are presented in Tab. 2 and Fig. 3. We find that the KPC-cF approach achieved adequately trained results in both subtasks for the actual korean data. The model pre-tuned with Kor-SemEval achieves the best performance in Aspect Category Detection (ACD).
Figure 3: Performance of ACD and ACP on KR3 test set during adaptive fine-tuning ($D_{T}$ and $D_{S}\rightarrow D_{T}$). Left: results with the addition of other fine-tuned BaselineXLM-R. th denotes the threshold for confidence of pseudo-labeling, and L denotes the threshold for filtering of LaBSE score; Right: BaselineXLM-R tuning compared in this paper. Blue line represents our proposed model, KPC-cF.
@@ 4 Discussion
In Phase 1, XLM-R, known for its proficiency in capturing cross-lingual representations, exhibits an underfitting tendency concerning the contextual disparities in aspect vocabulary within a single task. This can be attributed to data scarcity relative to model availability for each classifier or viewed as a limitation in single task using SPM in low-resource Korean ABSA Son et al. (2023). Nevertheless, in the NLI task, it showcases potential by outperforming mBERT, guided by "aspect". Conversely, mBERT demonstrates stable results in both single and NLI tasks, exhibiting an overall accuracy increase, particularly in the NLI task (Appendix D).
@@
Looking at Figure 4, In BaselineXLM-R+TR, the Target None class exhibits relatively high softmax probabilities, while data with polarity show relatively low probabilities, indicating significant confidence in polarity and asymmetric learning between classes. The results of KPC-cF demonstrate that additional fine-tuning with filtered data significantly increases the maximum probability of the model for Target Polarity. Simultaneously, with a reduction of over 40% in data, it prevents biased class learning, forming similar distributions among Target classes. In addition, by providing shallow features as input to a low-complexity LR model, we further confirmed the efficacy of the inherent characteristics of the filtered data through observed variations in training outcomes (see Tab. 3). Our framework suggests improving task stability and reliability by setting an optimal threshold to adjust biased classes or data with uncertain classification outcomes. Building upon Ouyang et al. Ouyang et al. (2023), we’ve shown that filtered training data is more effective for performance than simply having a large amount of information in real-world settings.
@@ 5 Conclusion
Aspect-Based Sentiment Analysis (ABSA) has been recognized as one of the most attractive subareas in text analytics and NLP. However, obtaining high-quality or ample-size labeled data has been one of the most essential issues hindering the development of ABSA. In this paper, we addressed the language gap in ABSA by building a pseudo-classifier. This involved aligning implicit features through dual filtering, further fine-tuning Korean NLI pairs with optimal pseudo-labels from a model trained on translated data. Additionally, we presented Kor-SemEval and KR3 train (pseudo labeled & filtered), testset (Gold) composed of Korean fine-grained set. We invite the community to extend Korean ABSA by providing new datasets, trained models, evaluation results, and metrics.
@@ B.1 Task description
In ABSA, Sun et al. Sun et al. (2019) set the task as equivalent to learning subtasks 3 (Aspect Category Detection) and subtask 4 (Aspect Category Polarity) of SemEval-2014 Task 4 simultaneously. A similar approach was adopted for Korean ABSA in restaurant reviews, aiming to develop a task-specific model through comparison of two PLMs, differing only in tokenization, vocabulary size, and model parameters Conneau et al. (2020). Defining a unified-serving model using multi-label-multi-class classification from a task-oriented perspective was considered impractical due to challenges in modifying pre-training set and the ongoing injection of additional data, rendering implicit mapping unattainable Sun et al. (2020); Ahmed et al. (2022); Qin and Joty (2022). Consequently, the problem has been redefined into two BERT-based classification tasks as outlined below.
@@ C.1 Dataset for Fine-Tuning and Test
Unlike the domains previously used for Korean sentiment classification Lee et al. (2020); Yang (2021); Ban (2022), Korean Restaurant Review with Ratings (KR3) is a restaurant review sentiment analysis dataset constructed through actual certified map reviews. In the case of restaurant reviews, words and expressions that evaluate positive and negative are mainly included, and real users often infer what a restaurant is like by looking at its reviews. Accordingly, Jung et al.3 constructed the KR3 dataset by crawling and preprocessing user reviews and star ratings of websites that collect restaurant information and ratings. KR3 has 388,111 positive and 70,910 negative, providing a total of 459,021 data plus 182,741 unclassified data, and distributed to Hugging Face [^4] https://huggingface.co/datasets/leey4n/KR3.
@@ C.2 Hyperparameter
In the Preliminary Experiment (Appendix D), we set the batch size to 3, this setting was applied to both single and NLI tasks. The max length was set to 512, and for epochs beyond 3, no significant performance improvement was observed, so the results from epoch 2 were noted. Subsequently, in Main Experiment (Sec. 3.1), we fine-tuned with a batch size of 16, following the pattern of previous experiments Karimi et al. (2021). However, no significant performance improvement was observed within 10 epochs, so we reported the results from epoch 4. Each reported metric is the average of runs with four different random seeds to mitigate the effects of random variation in the results.
@@ Appendix E Limitations and Future Work
Acknowledging limitations is crucial before delving into future work. While mBERT exhibits somewhat inadequate transfer effects, which can be attributed to its smaller model size. Due to pre-training solely on translation data, the scores are densely populated, resulting in limited performance improvement through additional training. Nevertheless, unlike jointly with Target None, the regression line of maximum probability for Target Polarity still trends upward according to LaBSE score (see Tab 1, Figure 5), indicating that ongoing evaluation of threshold points for the two scores can be done amidst more diverse and precise patterns. Meanwhile, we conducted direct inference on individual NLI inputs through a demo [^5] https://huggingface.co/KorABSA. Upon detecting various aspects within the reviews, we observed a slight predominance towards the "None" class within the class-wise similar probability distributions per sentence in TR training, while a more concentrated prediction tendency towards the binary class was observed in PL training. Furthermore, a tendency for one class to dominate over two or more aspect classes within review sentences was identified. This appears to stem from the class imbalance in TR and contextual differences Jin et al. (2024) within all review data (i.e., TR & PL).
@@
PL > L0.15 $\cap$ PL > Th0.5 (BaselineXLM-R) 37 0.243%
PL < L0.15 $\cap$ PL > Th0.5 (BaselineXLM-R) 68 0.446%
Table 7: Quality assessment results of PL filtering by two Baseline outputs; Target None ratio is 31.1% and 0.203% lower in the PL-cF threshold.
@@
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Tab. 7, 8). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None.
Furthermore, our threshold filtering that assigns a high-confidence Target None label for connotations with high similarity but lower errors supports a better distinction between semantic similarity and causal relationships. This is particularly true when excluding personalized terms and korean proper nouns like ‘Fixture (i.e., local native)’ used by reviewers. As a method to mitigate Spuriousness Haig (2003); Udomcharoenchaikit et al. (2022), we expect that our upper filtering will enable Target None predictions to function as data that helps the model learn the actual entailment relationships between sentences without relying solely on surface-level similarity in NLI pairs.

### v3

@@ Abstract
Investigations into Aspect-Based Sentiment Analysis (ABSA) for Korean industrial reviews are notably lacking in the existing literature. Our research proposes an intuitive and effective framework for ABSA in low-resource languages such as Korean. It optimizes prediction labels by integrating translated benchmark and unlabeled Korean data. Using a model fine-tuned on translated data, we pseudo-labeled the actual Korean NLI set. Subsequently, we applied LaBSE and MSP-based filtering to this pseudo-NLI set as implicit feature, enhancing Aspect Category Detection and Polarity determination through additional training. Incorporating dual filtering, this model bridged dataset gaps, achieving positive results in Korean ABSA with minimal resources. Through additional data injection pipelines, our approach aims to utilize high-resource data and construct effective models within communities, whether corporate or individual, in low-resource language countries. Compared to English ABSA, our framework showed an approximately 3% difference in F1 scores and accuracy. We release the dataset and our code for Korean ABSA, at this link [^1] https://github.com/namkibeom/KPC-cF.
@@ 2.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur and Turchi Balahur and Turchi (2012) has shown that Machine Translation (MT) systems can obtain training data for languages other than English in general sentiment classification. However, existing research using translation data for alignment and alignment-free methods Li et al. (2020); Zhang et al. (2021) inadequately address the challenge of universal knowledge transfer for linguistic subgroups like Korean. Also, although it was a different domain at Zhou et al. Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature $\Phi$ between translated source data $\bm{D_{S}}$ and actual target language data $\bm{D_{T}}$. Therefore, we attempted the following two phases to assess the impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Fig. 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, similar to the findings of Hoshino et al. (2024), the optimal baseline model for Korean reasoning is identified from the pool of models trained and assessed utilizing the translation dataset. In Phase 2, we fine-tune the baseline model $\Psi_{pre}(\Phi(D_{S});\theta_{D_{S}})$, which was effective in training on $D_{S}$, by additionally incorporating pseudo-labeled actual $D_{T}$. Employing the tuned model $\Psi_{post}(\Phi(D_{T});\theta_{{D_{S}}\rightarrow D_{T}})$, we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, LaBSE and confidence score filtering are performed to enhance implicit features $\Phi(D_{T})$. Detailed description of our Language Adaptation for aligned ABSA task is provided in Appendix§B.
@@
7:         $score_{MSP}\leftarrow$ $\textit{MSP}_{\Psi_{pre}}$($\hat{y}\mid x_{s},x_{a}$)
@@ 2.2 LaBSE based Filtering
We aimed to apply the LaBSE scoring to the actual Korean data $D_{T}$, KR3, intending to facilitate flexible learning compared to the translated data $D_{S}$, Kor-SemEval (see Appendix§C.1).
@@ 2.3 Confidence score Filtering
Meanwhile, we need to develop a classifier $\Psi_{post}$ capable of optimal predictions on the $D_{T}$ test, which can be considered as out-of-distribution data separate from the $D_{S}$. Drawing on previous research Arora et al. (2021), we expect that language shifts (i.e., $D_{S}\leftrightarrow D_{T}$) embody both Background and Semantic shift characteristics. To ensure robust learning in both aspect detection and sentiment classification, we introduce additional thresholding on Maximum Softmax Probability (MSP; Hendrycks and Gimpel 2017) after LaBSE-based filtering on the $D_{T}$ train set. When considering an input $x=(x_{s},x_{a})\in\mathcal{X}$ and its corresponding pseudo-label $\hat{y}\in\mathcal{Y}$, the score $s(x)$ for MSP is expressed as:
$s_{\textit{MSP}}(x)=\max_{k\in\mathcal{Y}}p_{\text{model}}(\hat{y}=k\mid x).$ (4)
@@
Table 2: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to total number (i.e., equivalent to Kor-SemEval) and capacity of samples from KR3 Train. The performances where filtered models outperform non-filtered ones under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Appendix§B.3.
@@ 3.1 Main Experiment: KR3 Test Set
Based on the results from Kor-SemEval, we observed analogous patterns between mBERT and XLM-RBase, notwithstanding their distinct properties (see Appendix§D). Accordingly, We opted for the Baseline-NLI approach (i.e., BaselinemBERT, XLM-R), which demonstrated the best performance, as the base model for Phase 2.
@@
Evaluation results on the KR3 test set are presented in Tab. 2 and Fig. 3. We find that the KPC-cF approach achieved adequately trained results in both subtasks for the actual korean data. The model pre-tuned with Kor-SemEval achieves the best performance in Aspect Category Detection (ACD).
Figure 3: Performance of ACD and ACP on KR3 test set during adaptive fine-tuning (e.g., $D_{T}$ and $D_{S}\rightarrow D_{T}$). Left: results with the addition of other fine-tuned BaselineXLM-R. th denotes the threshold for confidence of pseudo-labeling, and L denotes the threshold for filtering of LaBSE score; Right: BaselineXLM-R tuning compared in this paper. Blue line represents our proposed model, KPC-cF.
@@ 4 Discussion
In Phase 1, XLM-R, known for its proficiency in capturing cross-lingual representations, exhibits an underfitting tendency concerning the contextual disparities in aspect vocabulary within a single task. This can be attributed to data scarcity relative to model availability for each classifier or viewed as a limitation in single task using SPM in low-resource Korean ABSA Son et al. (2023). Nevertheless, in the NLI task, it showcases potential by outperforming mBERT, guided by "aspect". Conversely, mBERT demonstrates stable results in both single and NLI tasks, exhibiting an overall accuracy increase, particularly in the NLI task (Appendix§D).
@@
Model Task Type KorSTS KLUE-STS
Zero-shot Inference
[…]
BaselineXLM-R+TR $D_{S}$ 4.53 1.18
BaselineXLM-R+ PL $D_{T}$0.8 2.59
BaselineXLM-R+TR+PL $D_{S}\rightarrow D_{T}$ 3.26 3.29
KPC-cF $D_{S}\rightarrow D_{T}$ 4.98 1.20
Table 3: Comparisons of intermediate layer features on general Korean STS. We report Spearman’s $\rho$ (%). Higher is better. KLUE-STS discussed in Appendix§E.
Looking at Figure 4, In BaselineXLM-R+TR, the Target None class exhibits relatively high softmax probabilities, while data with polarity show relatively low probabilities, indicating significant confidence in polarity and asymmetric learning between classes. The results of KPC-cF demonstrate that additional fine-tuning with filtered data significantly increases the maximum probability of the model for Target Polarity. Simultaneously, with a reduction of over 40% in data, it prevents biased class learning, forming similar distributions among Target classes. Additionally, for polarity data where the target is detected, the regression line shows that higher LaBSE scores correspond to increased confidence. Enhancing this characteristic of the model, our KPC-cF demonstrates an improved ability to detect general semantic relationships in Korean, as shown in Tab. 3. In this experiment, we generated intermediate embeddings for KorSTS ham-etal-2020-kornli sentence pairs in a zero-shot manner and measured their similarity, presenting the results in terms of Spearman’s correlation ($\rho$) with the Ground Truth similarities. On the other hand, by providing word level features as input to a low-complexity LR model, we further confirmed the efficacy of the basic characteristics of the filtered data through observed variations in training outcomes (see Tab. 4). Our framework suggests improving task stability and reliability by setting an optimal threshold to adjust biased classes or data with uncertain classification outcomes.
@@
4.1 Further Analysis
To analyze the types of anticipatory effects that enable language models to exhibit improved performance in filtering corpora, we conducted a qualitative analysis by examining 116 cases where KPC-cF outputs matched Ground Truth but not those of the misaligned model (i.e., BaselineXLM-R+TR+PL). Interestingly, we identified the following four patterns that complement the predictive behavior of TR+PL futher training: (1) False Detection Refining: TR+PL incorrectly predicted aspects not mentioned in the reviews with a specific polarity, but KPC-cF refined these to ‘None’. (2) Mismatch Refining: TR+PL incorrectly predicted aspects mentioned in the reviews with a completely different polarity, which KPC-cF refined to the correct polarity. (3) Multi Context Detecting: KPC-cF detected contexts within a review with multiple aspects, correctly assigning the ‘None’ overlooked by TR+PL to the correct polarity. (4) Single Context Detecting: KPC-cF detected contexts within a single-aspect review and correctly assigned the polarity that TR+PL had overlooked as ‘None’.
Observed Instances %
(gt-matched, not $\hat{y}_{\text{TR+PL}}$)
False Detection Refining 34.38%
Mismatch Refining 31.25%
Multi  Context Detecting 18.75%
Single Context Detecting 15.62%
Table 5: Qualitative assessment results of KPC-cF predictions on the KR3 test set; We analyze aligned model outcomes with Refining and Detecting patterns.
Patterns 1 and 2 emphasize the role of KPC-cF in providing error feedback. The filtered PL corpus indicates the presence of more significant polarities, guiding responses from the model. In contrast, patterns 3 and 4 demonstrate that the filtering framework allows the language model to utilize aligned aspect information. This adaptation of cohesive contexts from the translation corpus optimizes problem-solving for the target task. The frequency of each pattern from our qualitative assessment is provided in Tab. 5. On the other hand, given the somewhat limited effectiveness compared to Refining, the Detecting patterns suggest that training initiated from TR still possesses both scalability limits and potential. Therefore, we propose additional analysis and directions in Appendix§E.
@@ 5 Conclusion
Aspect-Based Sentiment Analysis (ABSA) has been recognized as one of the most compelling subfields in text analytics and NLP. However, obtaining high quality labeled data has been one of the most important issues hindering the developments of ABSA. In this study, we addressed the language gap in ABSA by constructing a pseudo-classifier. This approach involved aligning implicit features through dual filtering and showcased its industrial and research potential via feature validation across various layers. Additionally, we presented Kor-SemEval, KR3 test (Gold) composed of Korean fine-grained set. We invite the community to extend Korean ABSA by providing new datasets, trained models, evaluation results, and metrics.
@@ B.1 Task description
In ABSA, Sun et al. Sun et al. (2019) set the task as equivalent to learning subtasks 3 (Aspect Category Detection) and subtask 4 (Aspect Category Polarity) of SemEval-2014 Task 4 simultaneously. A similar approach was adopted for Korean ABSA in food service reviews, aiming to develop a task-specific model through comparison of two PLMs, differing only in tokenization, vocabulary size, and model parameters Conneau et al. (2020). Defining a unified-serving model using multi-label-multi-class classification from a task-oriented perspective was considered impractical due to challenges in modifying pre-training set and the ongoing injection of additional data, rendering implicit mapping unattainable Sun et al. (2020); Ahmed et al. (2022); Qin and Joty (2022). Consequently, the problem has been redefined into two BERT-based classification tasks as outlined below.
@@ C.1 Dataset for Fine-Tuning and Test
Unlike the domains previously used for Korean sentiment classification Lee et al. (2020); Yang (2021); Ban (2022), Korean Restaurant Review with Ratings (KR3) is a comprehensive dataset encompassing various food service establishments, constructed from actual certified map reviews. In the case of restaurant reviews, words and expressions that evaluate positive and negative are mainly included, and real users often infer what a restaurant is like by looking at its reviews. Accordingly, Jung et al.3 constructed the KR3 dataset by crawling and preprocessing user reviews and star ratings of websites that collect broad food service information and ratings. KR3 has 388,111 positive and 70,910 negative, providing a total of 459,021 data plus 182,741 unclassified data, and distributed to Hugging Face [^4] https://huggingface.co/datasets/leey4n/KR3.
@@ C.2 Hyperparameter
In the Preliminary Experiment (Appendix§D), we set the batch size to 3, this setting was applied to both single and NLI tasks. The max length was set to 512, and for epochs beyond 3, no significant performance improvement was observed, so the results from epoch 2 were noted. Subsequently, in Main Experiment (Sec. §3.1), we fine-tuned with a batch size of 16, following the pattern of previous experiments Karimi et al. (2021). However, no significant performance improvement was observed within 10 epochs, so we reported the results from epoch 4. Each reported metric is the average of runs with four different random seeds to mitigate the effects of random variation in the results.
@@ Appendix E Limitations and Future Work
Acknowledging limitations is crucial before delving into future work. While mBERT exhibits somewhat inadequate transfer effects, which can be attributed to its smaller model size. Due to pre-training solely on translation data, the scores are densely populated, resulting in limited performance improvement through additional training. Nevertheless, unlike jointly with Target None, the regression line of maximum probability for Target Polarity still trends upward according to LaBSE score (see Tab 1, Figure 5), indicating that ongoing evaluation of threshold points for the two scores can be done amidst more diverse and precise patterns. Meanwhile, we conducted direct inference on individual NLI inputs through a demo [^5] https://huggingface.co/KorABSA. Upon detecting various aspects within the reviews, we observed a slight predominance towards the ‘None’ class within the class-wise similar probability distributions per sentence in TR training, while a more concentrated prediction tendency towards the binary class was observed in PL training. Furthermore, a tendency for one class to dominate over two or more aspect classes within review sentences was identified. This appears to stem from the class imbalance in TR and contextual differences Jin et al. (2024) within all review data (i.e., TR & PL).
@@
PL > L0.15 $\cap$ PL > Th0.5 (BaselineXLM-R) 2787 18.31%
PL < L0.15 $\cap$ PL > Th0.5 (BaselineXLM-R) 7470 49.06%
Table 9: Quality assessment results of PL filtering by two Baseline outputs; Target None ratio is 31.1% and 30.75% lower in the PL-cF threshold.
@@
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Tab. 9, 10). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None.
Model Task Type KLUE-STS-Mid.
Zero-shot Eval.
[…]
BaselineXLM-R+TR $D_{S}$9.0
BaselineXLM-R+ PL $D_{T}$ 3.24
BaselineXLM-R+TR+PL $D_{S}\rightarrow D_{T}$4.3
KPC-cF $D_{S}\rightarrow D_{T}$ 3.21
Table 11: Comparisons of intermediate layer features on KLUE-STS. We report Spearman’s $\rho$ (%). Higher is better. Data within the mid-range score were used.
Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. This limitation is evident not only in the qualitative analysis of specific data point errors but also in the KLUE-STS Park et al. (2021) results in Tab. 3. However, we have confirmed through various evaluations (Tab. 2, 4, 5, 10, 11 and Fig. 3) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations Haig (2003); Udomcharoenchaikit et al. (2022), we propose upper filtering that enables Target None predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs.